if a pipe flowing water has a cross sectional area nine times greater at point 2 than point 1, what would be the relation of flow of speed at the two points
Answer #1

Sol::

Answer is:
   "The flow speed at point 1 is nine times that at point 2"
Using the continuity equation we have
A1V1=A2V2
So, as per given question, the cross-sectional area at point 2 is nine times greater that at point 1 (i.e 9A1=A2).

So,from continuity equation
A1= 1 and A2 = 9.

=> 1*V1=9*V2

=> V1/V2=9/1
=> V1=9V2

Hence the Flow speed at point 1 is nine times that at point 2.


Hope will get RATED
THANK YOU